OpenStudy (anonymous):

Find x such that x + 4, 3x -9, and 2x +8 are consecutive terms of an arithmetic sequence.

OpenStudy (solomonzelman):

x is not a negative number, because if not, then `3x-9` has the greatest absolute value, and can't be the middle term. x is not zero, because 4,-9, 8 is not a sequence. So it must be a positive number. (this is what I got so far)

OpenStudy (solomonzelman):

x is big enough to make `3x-9` bigger than `x+4` so x + 4=3x -9 x +13=3x 2x=13 x=7.5 So x is bigger than 8. The 2x+8 is bigger than 3x-9 2x+8>3x-9 17>x So I am thinking it is 8<x<17 Okay then if it has to be an integer.... (idk why I am assuming that though) x=10 random guess.

OpenStudy (solomonzelman):

x + 4, 3x -9, 2x +8 10+4, 3(10)-9, 2(10)+8 14, 21, 28

OpenStudy (jdoe0001):

so we have 3 terms, and this is an arithmetic sequence meaning the next term is itself plus "something" so x + 4, 3x -9, 2x +8 so let's say that "something" is x so we could say that 3x - 9 = prior term + x 3x - 9 = (x+4) + x <--- solve for "x"

OpenStudy (anonymous):

So do I combine like terms on the right side?

OpenStudy (solomonzelman):

but then the x is not 10

OpenStudy (jdoe0001):

well... just solve or simplify for "x" :)

OpenStudy (anonymous):

Alright thank you both! c:

OpenStudy (jdoe0001):

yw

OpenStudy (xapproachesinfinity):

it is 10 the way jdoe001 did it was correct but he shouldn't put x as the common difference you need to put another variable there

OpenStudy (xapproachesinfinity):

let make is n (x+4)+n= the next term=3x-9 3x-9+n=2x+8

OpenStudy (xapproachesinfinity):

first solve for n and then solve for x ^_^

OpenStudy (jdoe0001):

hmmm thought about using "y" BUT that sorta will get a bit in the way of getting "x" methinks

OpenStudy (xapproachesinfinity):

but you that x is different variable than the common difference? no? so you need another variable y or call it whatever

OpenStudy (jdoe0001):

yes hhhhmmm.... ok

OpenStudy (jdoe0001):

well... I checked on the 1st two terms... and 13 checks out it doesn't for the last one though

OpenStudy (xapproachesinfinity):

n=2x-13 and n=-x+17

OpenStudy (xapproachesinfinity):

so x=10, nd 10 checks nicely

OpenStudy (jdoe0001):

ohhh I see..... d = d and then it'd be an equation based on "x"

OpenStudy (xapproachesinfinity):

Yes^_^

OpenStudy (jdoe0001):

@jherrera98 confused about what @xapproachesinfinity just said?

OpenStudy (anonymous):

Yes haha

OpenStudy (jdoe0001):

ok....lemme write it

OpenStudy (xapproachesinfinity):

what @SolomonZelman wasn't clear to me but somehow he got 10

OpenStudy (jdoe0001):

he did a rundown trial and error

OpenStudy (jdoe0001):

@jherrera98 one sec

OpenStudy (anonymous):

So solving for x isn't the answer?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (xapproachesinfinity):

but he used intervals how could he know that it is integer it could be any real?

OpenStudy (solomonzelman):

yes, I can't thing logically about thinking logically. I can use my brain, but unable to explain how I use it.....

OpenStudy (solomonzelman):

And I assumed it was an integer, because the question would have been very hard otherwise.

OpenStudy (anonymous):

but since it's a arithmetic sequence you have to be adding something don't you?

OpenStudy (xapproachesinfinity):

hehe, that's not a good way to do this it could really be a rational number hehe

OpenStudy (solomonzelman):

it can be an irrational number too

OpenStudy (xapproachesinfinity):

jherrera98 what you are adding is n=7

OpenStudy (xapproachesinfinity):

that would make way harder hehe

OpenStudy (anonymous):

What if I just equal them all to zero & find x?

OpenStudy (xapproachesinfinity):

i still need to observe this trial and error method hehe

OpenStudy (xapproachesinfinity):

No you can't. you are dealing with arithmetic sequence therefore you need to do it using the definition of an arithmetic sequence we know that for arithmetic sequence you add or subtract something to get the next term we call that common difference

OpenStudy (anonymous):

So there's many ways to solve this?

OpenStudy (anonymous):

I can't find d so idk what to do from there

OpenStudy (xapproachesinfinity):

Don't know! the obvious one is what jdoe did

OpenStudy (jdoe0001):

hmmmm not making much headway here either btw =)

OpenStudy (anonymous):

I'm even more lost now haha

OpenStudy (xapproachesinfinity):

the only thing i would thing of is that method! where do you lost?

OpenStudy (anonymous):

I'm lost once the common difference isn't the same for the last equation, 2x + 8

OpenStudy (jdoe0001):

see if you go this way \(\large { \begin{array}{cccllll} x+4&3x-9&2x+8\\ \hline\\ &(x+4)+d&(3x-9)+d \end{array} \\ \quad \\ \implies \begin{cases} 3x-9=(x+4)+d\to \frac{3x-9}{x+4}=d\\ 2x+8=(3x-9)+d\to \frac{2x+8}{3x-9}=d \end{cases} \\ \quad \\ d=d\implies \cfrac{3x-9}{x+4}=\cfrac{2x+8}{3x-9} }\) it ends up with \(\bf x^2-10x+7=0\ which .... is not very feasible

OpenStudy (jdoe0001):

ahemm with \(\bf x^2-10x+7=0\) that is, which is not very feasible either

OpenStudy (xapproachesinfinity):

No the common difference is indeed the same 14, 21,28

OpenStudy (jdoe0001):

ohhh man... I see my mistake now.. anyhow...lemme redo all that dohh =(

OpenStudy (anonymous):

So how do I find x now? .-.

OpenStudy (xapproachesinfinity):

you did multiplication lol,

OpenStudy (jdoe0001):

\(\large { \begin{array}{cccllll} x+4&3x-9&2x+8\\ &(x+4)+d&(3x-9)+d \end{array} \\ \quad \\ \implies \begin{cases} 3x-9=(x+4)+d\to 3x-9-(x+4)=d\\ 2x+8=(3x-9)+d\to 2x+8-(3x-9)=d \end{cases} \\ \quad \\ d=d\implies 3x-9-(x+4)=2x+8-(3x-9) }\) so.. there... now just a matter of solving for "x" :)

OpenStudy (jdoe0001):

see it @jherrera98 ? =)

OpenStudy (xapproachesinfinity):

we have solved already for x above, it think you are not following lol

OpenStudy (jdoe0001):

heheh k

OpenStudy (anonymous):

So x=10 is the final answer? haha

OpenStudy (xapproachesinfinity):

yes, what else? that's what the question is looking for you back to the terms and check you get 14,21,28 which true this sequence has 7 as the common difference

OpenStudy (jdoe0001):

\(\large { \begin{array}{cccllll} x+4&3x-9&2x+8\\ \hline\\ &(x+4)+d&(3x-9)+d \end{array} \\ \quad \\ \implies \begin{cases} 3x-9=(x+4)+d\to 3x-9-(x+4)&=d\\ 2x+8=(3x-9)+d\to 2x+8-(3x-9)&=d \end{cases} \\ \quad \\ d=d\implies 3x-9-(x+4)=2x+8-(3x-9) \\ \quad \\ 3x-9-x-4=2x+8-3x+9\implies 2x-13=-x+17 \\ \quad \\ 3x=30\implies x=\cfrac{\cancel{ 30 }}{\cancel{ 3 }}\to ? }\)

OpenStudy (anonymous):

Ok now I see it! Blonde moment lol Thank you all so much!

OpenStudy (xapproachesinfinity):

YW!